Price 100 200 250 350 Quantity How much tax revenue is raised? $800 $400 S1400 $700
Answer #1

Answer

The correct answer is (d) $700.

Suppose Government impose tax on Producers. This imposition of tax will result in increase in marginal cost of selling the good and hence at any given price they will supply less and thus supply will shift to the left.

We can see from above graph that Supply shifted from S to S'.

Equilibrium occurs at that quantity where Supply and demand intersect. We can see from above graph that New Supply(S') and Demand(D) intersects when Quantity = 100.

Thus quantity that are produced and consumed = 100.

We can also see from above that New Supply(S') and Demand(D) intersects when Price = 22. Hence Market price = 22 and thus price consumers are paying = 22.

Vertical distance between supply after tax and supply before tax = amount of tax = 22 - 15 = 7.

Thus From 22 that a producer receives per unit from buyers, they have paid $7 as tax and thus receives only 22 - 7 = 15 per unit.

Thus, Buyers paid price 22 per unit, sellers received 15 per unit and government received tax = 7 per unit.

Tax revenue = Per unit Tax*Quantity = 7*100 = 700.

Hence Government has raised Tax Revenue worth $700.

Hence, the correct answer is (d) $700.